# Experimental Observation of Motion of Ions in a Resonantly Driven Plasma Wakefield Accelerator

### v3

@@ Abstract
Figure 2a shows the time-resolved measured proton bunch density $n_{p^{+}}(x,\xi)$ with $N_{p^{+}}$=(2.8$\pm$0.1)$\times$1011 after propagation in vacuum (no plasma). The distribution is approximately bi-Gaussian with a transverse rms size of $\sigma_{\mathrm{x,y,SC}}=625$$\mu$m ($\gg\Delta y$). Figure 2 shows the density after propagating in xenon (b) and argon (c) plasmas with $n_{\mathrm{pe}}$=(4.8$\pm$0.2)$\times$1014 cm-3, the highest density reachable with helium. These show typical features of successful SMI PhysRevLett.122.054802 ; PhysRevLett.122.054801 : observable microbunch structure when using shorter time-windows (73 ps, Supplemental Material) PhysRevLett.122.054802 ; decrease of the transverse bunch size (visible from the front of the bunch to $\xi\simeq$ 250 ps), caused by adiabatic focusing of the bunch in plasma verra2022adiabatic ; signal decrease (visible for $\xi\lesssim 200$ ps, when compared to the no-plasma case on Fig. 2a), caused by the increase of proton divergence along the bunch EAAC2023SC . The divergence increases because the transverse wakefield amplitude increases due to resonant wakefield excitation. Protons with larger transverse momentum diverge more during vacuum propagation downstream of the plasma exit, leading to a lower bunch density measured with the streak camera because of the effect of the slit. Very little to no signal is observed for $\xi\lesssim$100 ps on Figs. 2b,c.
@@
The bunch density measured with helium (Fig. 2d) closely resembles those with argon (Fig. 2c) or xenon (Fig. 2b) from the front of the bunch to $\xi\sim-100$ ps, showing essentially the same SMI development and wakefield growth along the bunch and plasma. Microbunches are visible with all three plasmas on shorter time-windows also in this case (Supplemental Material). However with helium (Fig. 2d) and for $\xi\lesssim-100$ ps, the bunch density increases again, leading to the appearance of a bunch tail, not present on Figs. 2b and c  [^3] The intensity of the bunch tail around $\xi\sim-300$ ps is higher than when no plasma is present, due to adiabatic focusing PhysRevLett.129.024802 ; verra2022adiabatic .. This is the signature expected from the effect of the motion of ions on self-modulation PhysRevLett.109.145005 ; 10.1063/1.4876620 . This occurs when the mass of the ions is sufficiently low and the amplitude of the wakefields is sufficiently high for motion of ions to suppress the wakefields in the back of the bunch. Where wakefields are suppressed, the SMI development is suppressed and protons acquire much less transverse momentum, leading to much smaller divergence and increased proton bunch density, visible as a tail on the time-resolved image (Fig. 2d). The expected inverse dependence with $m_{i}$ is confirmed, as, with this amplitude of wakefields, only the bunch density measured with the lightest ions (helium) is disturbed. Simultaneous measurements with two streak cameras with orthogonal slits recording $n_{p^{+}}(x,\xi)$ and $n_{p^{+}}(y,\xi)$ (Supplemental Material) confirm that the core of the bunch and its tail are radially symmetric, as expected.
@@
Figure 4: Numerical simulation results (average of five simulations with different noise in the bunch particle distribution) with experimental parameters from Fig. 2 ($n_{pe}=4.8\times 10^{14}$ cm-3 and $N_{p^{+}}=2.8\times 10^{11}$) . Figures a-d: bunch distributions at the location of the streak camera screen and with: immobile (a), xenon (b), argon (c), and helium (d) ions. Effect of the streak-camera slit included. Figures e-h: corresponding densities of plasma ions at the plasma exit ($z$=10 m). Orange lines: envelope of the average transverse wakefields ($W_{\mathrm{r}}$) within $c/\omega_{pe}$, also at $z$=10 m. Error-bars: standard deviation of the simulation results. Numerical simulation result with argon and $n_{pe}=9.3\times 10^{14}$ cm-3 ($2n_{pe}$) in Supplemental Material.
@@
Numerical simulation results also show that the time for the tail to form along the bunch scales with $m_{i}^{-1/3}$ (Supplemental Material). This scaling is logical, as the tail formation is due to the loss of coherence in plasma electron motion or wavebreaking, which has been demonstrated to scale with $m_{i}^{-1/3}$ in Ref. lotovion . All experimental results obtained are consistent with this scaling.

### v4

@@ Abstract
Figure 2a shows the time-resolved measured proton bunch density $n_{p^{+}}(x,\xi)$ with $N_{p^{+}}$=(2.8$\pm$0.1)$\times$1011 after propagation in vacuum (no plasma). The distribution is approximately bi-Gaussian with a transverse rms size of $\sigma_{\mathrm{x,y,SC}}=625$$\mu$m ($\gg\Delta y$). Figure 2 shows the density after propagating in xenon (b) and argon (c) plasmas with $n_{\mathrm{pe}}$=(4.8$\pm$0.2)$\times$1014 cm-3, the highest density reachable with helium. These show typical features of successful SMI PhysRevLett.122.054802 ; PhysRevLett.122.054801 : observable microbunch structure when using shorter time-windows (73 ps, Supplemental Material SupplementalMaterial PhysRevLett.122.054802 ; decrease of the transverse bunch size (visible from the front of the bunch to $\xi\simeq$ 250 ps), caused by adiabatic focusing of the bunch in plasma verra2022adiabatic ; signal decrease (visible for $\xi\lesssim 200$ ps, when compared to the no-plasma case on Fig. 2a), caused by the increase of proton divergence along the bunch EAAC2023SC . The divergence increases because the transverse wakefield amplitude increases due to resonant wakefield excitation. Protons with larger transverse momentum diverge more during vacuum propagation downstream of the plasma exit, leading to a lower bunch density measured with the streak camera because of the effect of the slit. Very little to no signal is observed for $\xi\lesssim$100 ps on Figs. 2b,c.
@@
The bunch density measured with helium (Fig. 2d) closely resembles those with argon (Fig. 2c) or xenon (Fig. 2b) from the front of the bunch to $\xi\sim-100$ ps, showing essentially the same SMI development and wakefield growth along the bunch and plasma. Microbunches are visible with all three plasmas on shorter time-windows also in this case (Supplemental Material SupplementalMaterial ). However with helium (Fig. 2d) and for $\xi\lesssim-100$ ps, the bunch density increases again, leading to the appearance of a bunch tail, not present on Figs. 2b and c  [^3] The intensity of the bunch tail around $\xi\sim-300$ ps is higher than when no plasma is present, due to adiabatic focusing PhysRevLett.129.024802 ; verra2022adiabatic .. This is the signature expected from the effect of the motion of ions on self-modulation PhysRevLett.109.145005 ; 10.1063/1.4876620 . This occurs when the mass of the ions is sufficiently low and the amplitude of the wakefields is sufficiently high for motion of ions to suppress the wakefields in the back of the bunch. Where wakefields are suppressed, the SMI development is suppressed and protons acquire much less transverse momentum, leading to much smaller divergence and increased proton bunch density, visible as a tail on the time-resolved image (Fig. 2d). The expected inverse dependence with $m_{i}$ is confirmed, as, with this amplitude of wakefields, only the bunch density measured with the lightest ions (helium) is disturbed. Simultaneous measurements with two streak cameras with orthogonal slits recording $n_{p^{+}}(x,\xi)$ and $n_{p^{+}}(y,\xi)$ (Supplemental Material SupplementalMaterial ) confirm that the core of the bunch and its tail are radially symmetric, as expected.
@@
Figure 4: Numerical simulation results (average of five simulations with different noise in the bunch particle distribution) with experimental parameters from Fig. 2 ($n_{pe}=4.8\times 10^{14}$ cm-3 and $N_{p^{+}}=2.8\times 10^{11}$) . Figures a-d: bunch distributions at the location of the streak camera screen and with: immobile (a), xenon (b), argon (c), and helium (d) ions. Effect of the streak-camera slit included. Figures e-h: corresponding densities of plasma ions at the plasma exit ($z$=10 m). Orange lines: envelope of the average transverse wakefields ($W_{\mathrm{r}}$) within $c/\omega_{pe}$, also at $z$=10 m. Error-bars: standard deviation of the simulation results. Numerical simulation result with argon and $n_{pe}=9.3\times 10^{14}$ cm-3 ($2n_{pe}$) in Supplemental Material SupplementalMaterial .
@@
Numerical simulation results also show that the time for the tail to form along the bunch scales with $m_{i}^{-1/3}$ (Supplemental Material SupplementalMaterial ). This scaling is logical, as the tail formation is due to the loss of coherence in plasma electron motion or wavebreaking, which has been demonstrated to scale with $m_{i}^{-1/3}$ in Ref. lotovion . All experimental results obtained are consistent with this scaling.